# and quantum information science

### v1

@@ Abstract
Quantum complexity quantifies the difficulty of preparing a state, or implementing a unitary, using limited resources. Over the past decade, quantum-complexity research has surged. Applications range from quantum computation to condensed matter and quantum gravity. Different communities apply different tools to quantum complexity, as well as define complexity differently. This review bridges the disciplines’ approaches. We cover multiple definitions of complexity, as well as their key properties and applications. In quantum information, random quantum circuits proved to be an invaluable tool in extracting properties of the complexity growth of discrete quantum circuits. In quantum many-body systems and quantum field theory (QFT), complexity is defined geometrically through geodesics. Complexity can be defined also in terms of state or operator spreading and tensor-network concepts. In addition to reviewing these approaches, we outline applications to simple quantum systems, quantum many-body models and QFTs, including conformal field theories (CFTs). Finally, we detail the relationship between complexity and gravitational observables within the anti-de Sitter-space (AdS)/CFT correspondence. This review thereby integrates recent quantum-complexity advances across fields.
@@ 1 Preface
This review’s main goal is to facilitate new, and consolidate existing, bridges between communities whose members study quantum complexity. The communities use different conceptual approaches, mathematical tools, and levels of rigor. Other reviews cover some of the topics discussed below [1, 2, 3]. However, we offer two novel perspectives. First, our diverse backgrounds enable us to cover extensive ground at the intersection of gravity, quantum information theory, quantum field theory (QFT), and quantum many-body systems. Second, this review reflects the state of the art in this fast-moving field in early 2025.
We begin with a historical overview. Quantum complexity originated in the field of quantum computation. The concept served as a mathematical tool in the optimization of unitary transformations’ implementations. This optimization has concerned users of early quantum computers, which have limited resources. Since 2009, tensor networks have bridged quantum computation to holographic studies [4]. Tensor networks are graphical structures, similar to quantum circuits, that can efficiently represent certain quantum states classically. Holography is a duality between gravitational physics in a bulk geometry and a quantum system on the geometry’s boundary. Tensor networks helped establish quantum complexity’s relevance to gravitational physics. Via the holographic correspondence, quantum-complexity studies expanded into QFT and the Sachdev-Ye-Kitaev (SYK) model. In parallel, researchers developed complexity-related quantities for characterizing operator growth (how operators evolve in the Heisenberg picture) under many-body quantum chaos. Finally, starting in 2021, physicists have proven several conjectured properties of complexity measures, enabling a more precise formulation of the holographic correspondence.
@@
Figure 1 depicts the review’s structure. A reader might traverse it via any of (at least) four paths, indicated by arrows in the figure. A common thread unites each path’s components. All paths begin in Sec. 2, a glossary for key concepts. A reader may study this section more or less thoroughly, depending on their background, and return to the section later as necessary. Let us now outline the paths consecutively.
We start with the leftmost arrow in Fig. 1. It points to Section 3, a dialogue inspired by Galilei’s famous discourse on cosmology. Our dialogue resembles a discussion at a conference: not all terms are sharply defined, and different speakers present different perspectives. Later sections provide greater rigor and more detail. The dialogue terminates the first, briefest path through the review.
The second path focuses on quantum-information-theoretic features of quantum complexity (see the second-leftmost sequence of arrows in Fig. 1). This path includes the definition of circuit complexity in Sec. 4. Section 9 presents applications of complexity to quantum information theory and many-body physics.
The third path’s protagonist is the evolution of complexity under Hamiltonian dynamics (see the second-rightmost sequence of arrows in Fig. 1). Section 4 describes how complexity grows under Hamiltonian evolution, using counting arguments, and under random circuits. The complexity’s time evolution inspired the definitions of geometrical observables in gravitational physics, as discussed in Sec. 7.
The right-hand side of Fig. 1 illustrates the fourth path through the review. It focuses on quantum-complexity definitions suited to continuous-time evolution. Sections 5, 6 and 8 introduce Nielsen’s complexity, the Krylov complexity, and complexity quantities inspired by tensor networks. All these measures elucidate thermalization, chaos, quantum-circuit optimization, and more.
Apart from the first path, all paths end in Sec. 10. There, we summarize the main results and open problems. Like Sec. 3, Sec. 10 contains a dialogue, for accessibility and the reader’s amusement. A contains a guide to the acronyms in this paper.
Some paths intersect at certain topics. For example, the holographic duality proposes that each quantum-complexity measure has an analogue in a gravitational system: a geometric observable. This intertwining surfaces in Sections 58.
@@ 2 Key notions
This review covers developments across disciplines, from high-energy physics and gravity through many-body physics to quantum information science. These disciplines feature different vocabularies. Here, we survey key notions from across the disciplines, to assist readers from all backgrounds with our multifaceted material.
@@
    AdS/CFT correspondence: A duality between (i) gravity in asymptotically anti-de Sitter (AdS) spacetime and (ii) a conformal field theory (CFT), typically a gauge theory, on the spacetime’s timelike boundary.
@@
    Black hole: A region of spacetime where gravity is so strong that no matter or light can escape. Mathematically, a black hole is represented by a solution to Einstein’s equations (or to another gravitational theory’s equations of motion). A black hole contains a singularity, where certain scalar functions of the curvature diverge.
@@
    Duality: An equivalence between two seemingly distinct descriptions of the same physics. The two descriptions related by a duality are often called sides.
@@
    Entanglement: A nonlocal relationship achievable between quantum particles. The relationship can lead to correlations (between outcomes of measurements of the particles) stronger than any achievable classically. An entangled state stores information inaccessible via local measurements. The entanglement entropy, among other measures, quantifies the entanglement between two systems.
@@
    Event horizon: A null hypersurface beyond which no matter or light can escape.
@@
    Fast scrambler: A quantum system that quickly distributes quantum information across all of its degrees of freedom. Generic out-of-time-ordered correlators (OTOCs, see the corresponding key notion) of a fast scrambler decay exponentially quickly, with the greatest possible exponent, during a certain time frame. Let $N$ denote the number of degrees of freedom; and $t$, time. A generic (possibly regularized) OTOC of a fast scrambler scales as
@@
    The Lyapunov-like exponent $\lambda_{\rm OTOC}$ is well-defined only in a semiclassical regime, where one can expand the OTOC in a small parameter ($\hbar$ or $1/N$, for instance). At the scrambling time $t_{*}\sim\lambda_{\rm OTOC}^{-1}\log(N)$, the first two terms in (1) equal each other: information about a local perturbation [$W$ in Eq. (3)] has scrambled throughout the system; $W(t)$ and $V$ have an order-1 commutator. A system’s scrambling time is as low as possible if the Lyapunov-like exponent is as large as possible: $\lambda_{\rm OTOC}=2\pi/\beta$ [5]. Systems that achieve this condition are called fast scramblers. Black holes and the large-$N$ SYK model are believed to be examples.
@@
    Holography: A duality stating that gravitational physics in a bulk region is equivalent to a quantum theory on the region’s boundary. Often, a holographer aims to identify a map between a bulk quantity and a boundary quantity. Identifying such a map, one matches the two quantities.
@@
    Jackiw-Teitelboim (JT) gravity: A two-dimensional gravitational theory with a coupling to a dilaton (a real scalar field $\Phi$[6]. Let $G_{\rm N}$ denote Newton’s constant; $\mathcal{M}$, a $d$-dimensional smooth manifold; $g$, the metric’s determinant; $R$, the Ricci scalar; $L$, the AdS radius; $K$, the extrinsic curvature of the induced metric on the manifold’s boundary, $\partial\mathcal{M}$; and $h$, the determinant of that induced metric. The JT-gravity action, excluding topological contributions, has the form
    $S_{\rm JT}[g,\Phi]\coloneqq\frac{1}{16\pi G_{\rm N}}\int_{\mathcal{M}}d^{d}x\,\sqrt{-g}\>\Phi\left(R+\frac{2}{L^{2}}\right)+\frac{1}{8\pi G_{\rm N}}\int_{\partial\mathcal{M}}d^{d-1}x\,\sqrt{-h}\>\Phi\left(K-\frac{1}{L}\right)\,.$ (2)
    The AdS/CFT duality interrelates (i) JT gravity in asymptotically AdS spacetime and (ii) the SYK model’s low-energy regime.
@@
    Kolmogorov complexity: The length of the shortest computer program that produces a given output.
@@
    Out-of-time-order correlator (OTOC): A correlation function intended to measure the spread of many-body entanglement across, and to signal chaos of, quantum systems. Let $S$ denote a quantum many-body system. Let $W$ and $V$ denote local operators that are unitary or Hermitian, depending on the context, and that act on far-apart subsystems of $S$. $W(t)$ follows from evolving $W$, in the Heisenberg picture, under the system’s dynamics. Let $\rho$ denote a state of $S$. The OTOC has the form [5]111 The OTOC often diverges in QFT. One therefore regularizes the OTOC: one removes the $\rho$ from the trace’s argument. Instead, most commonly, one places a $\rho^{1/4}$ after each remaining factor. Alternative regularization prescriptions differ in the power of $\rho$ and in the operators’ ordering.
    $F(t)={\rm Tr}\left(W^{\dagger}(t)\,V^{\dagger}\,W(t)\,V\,\rho\right)\,.$ (3)
    The OTOC quantifies the failure of $W(t)$ to commute with $V$. Relatedly, the OTOC quantifies how much a local perturbation $V$ affects a far-away operator $W$ later. Often, when referring to an OTOC, a text actually refers implicitly to the OTOC’s real part.
@@
    Quantum circuit: A sequence of quantum gates. In a conventional circuit diagram, time runs from left to right. Time is discretized into slices, or layers. Multiple gates can act during each layer (if they commute with each other and if, at most, one gate acts on each qubit). The number of layers is the circuit’s depth.
@@
    Quantum complexity: The least number of quantum logic gates, selected from a chosen set, in any circuit that (i) implements a target unitary or (ii) prepares a target state from a simple initial state. High-energy theorists sometimes call quantum complexity quantum computational complexity, which can have a different meaning in computer science. In sections that provide explicit analytic bounds, we will emphasize the difference between the exact circuit complexity $\mathcal{C}_{0}(U)$ of a unitary $U$, measuring the minimal number of gates to effect $U$ exactly, and the approximate circuit complexity $\mathcal{C}_{\delta}(U)$, the minimal number of gates to approximate $U$ up to an error of $\delta$. When this distinction is not central to the material presented, we drop the subscript.
@@
    Quantum gate: An elementary unitary operation that acts on a quantum system, usually a set of qubits. A $k$-local quantum gate acts on just $k$ qubits. Example 1-qubit gates include the Hadamard gates; and example 2-qubit gates include the controlled-NOT, or CNOT, gate [10].
@@
    Random state: A quantum state selected from a Hilbert space, typically according to the Haar measure, the unique unitarily invariant measure. One can think of a (Haar-)random state as “uniformly randomly” selected.
@@
    The SYK model admits simplifications when taking $N\rightarrow\infty$ and $q\rightarrow\infty$ keeping the ratio $\lambda_{\text{SYK}}=2q^{2}/N$ fixed, which is often referred to as double scaled SYK (DSSYK).
@@
    Tensor networks: A tensor network is a graphical representation of a quantum state [11]. The network consists of basic units, called tensors, that are multilinear maps. A tensor’s inputs and outputs, labeled by indices, represent quantum degrees of freedom. To evaluate a tensor network, one contracts the tensors, summing over indices. Tensor networks are useful, e.g., in variational studies of ground states. Also, using tensor networks, one can bound entanglement entropies in a simple, graphical manner.
    An important type of tensor network is the multiscale entanglement renormalization ansatz (MERA) [12]. MERA efficiently approximates scale-invariant quantum systems’ ground states. A continuous version (cMERA) has been proposed to approximate free CFTs’ ground states well [13].
@@
    Thermofield-double state: Consider a quantum system governed by a Hamiltonian $H$ with eigenvalues $E_{n}$ and eigenstates $|n\rangle$. Denote by $Z_{\beta}$ the partition function at an inverse temperature $\beta$. Two copies of the system can jointly be in the thermofield-double state
@@
    Each copy is in a finite-temperature canonical (thermal) reduced state.
@@
    Universal gate set: A set of quantum gates such that, using gates drawn from the set, one can approximately effect every unitary operator arbitrarily precisely.

### v2

@@ Abstract
Quantum complexity quantifies the difficulty of preparing a state or implementing a unitary transformation with limited resources. Applications range from quantum computation to condensed matter physics and quantum gravity. We seek to bridge the approaches of these fields, which define and study complexity using different frameworks and tools. We describe several definitions of complexity, along with their key properties. In quantum information theory, we focus on complexity growth in random quantum circuits. In quantum many-body systems and quantum field theory (QFT), we discuss a geometric definition of complexity in terms of geodesics on the unitary group. In dynamical systems, we explore a definition of complexity in terms of state or operator spreading, as well as concepts from tensor-networks. We also outline applications to simple quantum systems, quantum many-body models, and QFTs including conformal field theories (CFTs). Finally, we explain the proposed relationship between complexity and gravitational observables within the holographic anti-de Sitter (AdS)/CFT correspondence.
Preprint number: YITP-25-39
@@ 1 Preface
The goal of this review is to facilitate the construction of new bridges between different fields of physics that study dynamical phenomena through the lens of quantum complexity. These fields define and use different notions of complexity, which require distinct methods and tools for analysis. Many of these developments are individually reviewed in [1, 2, 3]. We hope to offer a complementary, integrative perspective at the intersection of gravity, quantum information theory, quantum field theory (QFT), and quantum many-body systems. This is a fast moving field and our review captures the state of the art in early 2025.
Historically, the study of quantum complexity originated in the field of quantum computation, generalizing ideas from classical computation. The basic question was, “How difficult is it to solve instances of a particular problem class?” Here, difficulty was measured in terms of a scarce resource like the minimal amount of time, space, or number of circuit gates necessary to implement an algorithm to solve problems exactly, approximately, or even just with high probability. Defined in this way, the concept of complexity functioned as a mathematical tool guiding the design of optimal problem-solving algorithms or circuits, or in the language of quantum mechanics, unitary transformations. Given the arbitrariness inherent in any computational model, the output of complexity theory was usually given in the form of bounds on the minimal amount of resource required to solve general instances of the problem class, as a function of some measure of the problem size. Many related questions have been asked in quantum information theory. For example, how many local operations are required to build a given entangled state of $N$ spins starting with a product state? The answer will depend quantitatively on the “amount” and “kind” of entanglement in the target state, both quantified by information theoretic measures. Local operations are privileged in such analyses because physical processes act locally, and nonlocal coordination or structures must be constructed from local ones. Many of these notions are reviewed in [4].
Over the last two decades, these sorts of questions have also turned out to be relevant for quantum gravity. This is because, in the “holographic” duality between $d+1$-dimensional quantum gravity in Anti-de Sitter (AdS) spacetimes and $d$-dimensional Conformal Field Theories (CFTs), entanglement patterns in the CFT state are geometrically encoded in the dual gravity theory [5]. Matrix Product and tensor network constructions of quantum states provide cartoon models for visualizing the holographic encoding of quantum information, with coarser structure built up deeper in the network and finer structure at the periphery. Questions about the structure, dynamical construction, and time evolution of entangled quantum states can then also be geometrized. We will discuss aspects of the rich literature on this subject as appropriate.
Within this holographic quantum gravity setting, important impetus for thinking about “complexity” as a physical quantity came from the observation that some aspects of geometry on the gravity side of the AdS/CFT duality could not possibly be related to quantum information or entanglement-related quantities in the CFT [6]. Specifically, it turns out that the volumes of spatial sections of black hole spacetimes keep increasing through to times exponential in the black hole entropy, far after the timescale in the dual theory where all entanglement related quantities have equilibrated. Now, if we think of time evolution in a physical theory as a quantum computation enacted by a local Hamiltonian, then the complexity of that computation, understood for example as the size of the minimal local circuit that could perform it, can increase for exponential time. This suggested the idea that some notion of quantum complexity may be geometrized as the volume of spatial section in the AdS gravities dual to CFTs – see Chapter 6 of the review [4] – and seeded many investigations of alternative notions of complexity in physical systems. The resulting ideas have been applied to an enormous range of problems ranging from thermalization in bounded systems, to quantum information in many-body physics, to quantum chaos and integrability, and to the quantization of gravity. Some of these ideas are also shedding light on foundational problems such as the origin of the entropy of black hole horizons, and whether information is lost during black hole evaporation. Some solutions to these problems suggest that black holes appear to have a horizon and to destroy information simply because the fine structure required to decode the information is inaccessible to simple probes (older approaches are discussed in [7] and new developments are reviewed in [8]). The new tools that are now available to define and measure quantum complexity may enable us to study such ideas quantitatively.
@@
Figure 1 depicts the review’s structure; the arrows indicate possible paths through the material. All paths begin in Sec. 2, a glossary for key concepts. A reader may study this section more or less thoroughly, depending on their background, and return to the section later as necessary. Let us now outline the paths consecutively.
The leftmost arrow in Fig. 1 points to Sec. 3, a dialogue inspired by Galilei’s famous discourse on cosmology. The dialogue is modeled on discussions at a conference: not all terms are sharply defined, and different speakers present different perspectives. Later sections provide greater rigor and detail. The dialogue terminates the first, briefest path through the review.
The second path, marked by the second-leftmost sequence of arrows in Fig. 1, focuses on quantum information theoretic features of complexity. This path defines circuit complexity in Sec. 4. Section 9 presents applications of complexity to quantum information theory and many-body physics.
The third path, marked by the second-rightmost sequence of arrows in Fig. 1, considers time evolution of complexity. Section 4 uses counting arguments to analyze the effects of Hamiltonian dynamics and of random circuits transforming a state. Some of this material has a bearing on proposed geometric duals of quantum complexity in the correspondence between quantum theories of gravity in asymptotically AdS spacetimes and Conformal Field Theories, as discussed in Sec. 7.
The right-hand side of Fig. 1 illustrates the fourth path through the review. It focuses on quantum complexity definitions that are suited to continuous-time evolution. Section 5 introduces the idea that complexity of time evolution can be measured in terms of lengths of geodesic paths on the unitary group from the identity to the time evolution operator. Sec. 6 introduces approaches that measure complexity of dynamics in terms of the spread of operators or states over time. Sec. 8 introduces approaches inspired by tensor networks.
Some of the main results and open problems are summarized in another dialogue in Sec. 10. A contains a guide to acronyms used in this review.
Many ideas about complexity in physical systems that are being studied today have antecedents in the classic works of Landauer, Bennet, Zurek and other pioneers of the application of information and computation theory to physics. Interested readers can learn more in the collection [9] where questions like “What is complexity?”, “What causes its increase?”, and “Is there a limit to its increase?” are discussed along with topics of current interest like the onset of chaos and the formation of complex states from simple initial conditions.
@@ 2 Key notions
Below we survey vocabularies from high-energy physics and quantum gravity through many-body physics and quantum information science to assist readers from different backgrounds.
@@
    AdS/CFT correspondence: A duality between gravity in asymptotically anti-de Sitter (AdS) spacetime and a conformal field theory (CFT) on the spacetime boundary.
@@
    Black hole: A spacetime region bounded by a horizon (see below) that conceals a singularity where certain scalar functions of the curvature diverge. Mathematically, black holes are classical solutions to Einstein’s equations or to another theory of gravity.
@@
    Duality: An equivalence between two seemingly distinct descriptions of the same physics.
@@
    Entanglement: Two degrees of freedom in a pure quantum state are entangled if their wavefunction does not factorize. If they are in a mixed state, they are considered entangled if their density matrix is not separable. This relationship can lead to the measurement of correlations stronger than any achievable classically. Entangled states store information inaccessible via measurements on individual parties. The von Neumann or entanglement entropy quantifies entanglement between systems.
@@
    Event horizon: A null hypersurface which bounds a region of trapped surfaces. Matter and light cannot escape to infinity through the event horizon.
@@
    Fast scrambler: A quantum system that quickly distributes information across degrees of freedom. Generic OTOCs (see below) of fast scramblers decay rapidly at early times:
@@
    where $N$ is the number of degrees of freedom and $t$ is time. $\lambda_{\rm OTOC}$ is defined via semiclassical expansion in a small parameter like $\hbar$ or $1/N$. At the scrambling time $t_{*}\sim\lambda_{\rm OTOC}^{-1}\log(N)$, the first two terms in (1) are equal. Black holes and the large-$N$ SYK model are believed to be examples of fast scramblers, with the largest possible $\lambda_{\rm OTOC}=2\pi/\beta$ [10].
@@
    Holography: A duality stating that gravitational physics in a bulk region is equivalent to a quantum theory on the region’s boundary. A holographic map between bulk and boundary quantities identifies bulk quantities with the dual boundary description.
@@
    Jackiw-Teitelboim (JT) gravity: A two-dimensional gravitational theory coupled to a dilaton (a real scalar field $\Phi$[11]. Let $G_{\rm N}$ be Newton’s constant; $\mathcal{M}$, a $2$-dimensional smooth manifold; $g$, the metric’s determinant; $R$, the Ricci scalar; $L$, the AdS radius; $K$, the extrinsic curvature of the induced metric on the manifold’s boundary, $\partial\mathcal{M}$; and $h$, the determinant of that induced metric. The JT-gravity action, excluding topological contributions, is:
[…]
    AdS/CFT duality relates JT gravity and the low-energy SYK model.
@@
    Kolmogorov complexity: The length of the shortest program producing a given output.
@@
    Out-of-time-order correlator (OTOC): A correlation function measuring the spread of many-body entanglement. Let $S$ be a quantum many-body system, and let $W$ and $V$ denote local unitary or Hermitian operators that act on distant subsystems of $S$. Let $W(t)$ be the Heisenberg-picture time-evolved operator $W$ and let $\rho$ denote a state of $S$. The OTOC [10]111 The OTOC often diverges in QFT. One therefore regularizes it by removing the $\rho$ from the trace’s argument. Instead, most commonly, one places a $\rho^{1/4}$ after each remaining factor. Alternative regularization prescriptions differ in the power of $\rho$ and in operator ordering.
    $F(t)={\rm Tr}\left(W^{\dagger}(t)\,V^{\dagger}\,W(t)\,V\,\rho\right)\,$ (3)
    quantifies the failure of $W(t)$ to commute with $V$, or how much a local perturbation $V$ later affects a far-away operator $W$. Texts often refer implicitly to the real part of the OTOC.
@@
    Quantum circuit: A sequence of quantum gates, often represented with time, discretized into layers, running from left to right. Multiple gates can act in each layer (if they commute and if at most one gate acts on each qubit). The number of layers is the circuit’s depth.
@@
    Quantum complexity: The least number of quantum logic gates, selected from a chosen set required to (i) implement a target unitary transformation or (ii) prepare a target state from a given initial one. High-energy theorists sometimes call this quantity quantum computational complexity, which can have a different meaning in computer science. In sections providing  analytic bounds, we will emphasize differences between the exact circuit complexity $\mathcal{C}_{0}(U)$ of a unitary $U$, and the approximate circuit complexity $\mathcal{C}_{\delta}(U)$, the minimal number of gates to approximate $U$ up to error $\delta$. When this distinction is unimportant, we drop the subscript.
@@
    Quantum gate: An elementary unitary operation that acts on a quantum system, usually a set of qubits. A $k$-local quantum gate acts on just $k$ qubits. Example 1-qubit gates include the Hadamard gate; and example 2-qubit gates include the controlled-NOT, or CNOT, gate [15].
@@
    Random state: A state selected from a Hilbert space according to the Haar measure. This measure is unitarily invariant; so (Haar-)random states are “uniformly randomly”.
@@
    The SYK model admits simplifications if we take $N\rightarrow\infty$ and $q\rightarrow\infty$ with the ratio $\lambda_{\text{SYK}}=2q^{2}/N$ fixed. This limit is often referred to as double scaled SYK (DSSYK).
@@
    Tensor networks: A graphical representation of quantum states [16] built from multilinear maps called tensors. A tensor’s inputs and outputs, labeled by indices, represent degrees of freedom. Tensor networks are evaluated by contracting tensors by summing over indices. Tensor networks assist variational studies of ground states, and can be used to bound entanglement entropies in a simple, graphical manner. One example, the multiscale entanglement renormalization ansatz (MERA) [17], efficiently approximates ground states of scale-invariant systems. A continuous version (cMERA) approximates ground states of free CFTs [18].
@@
    Thermofield-double state: Consider a quantum system with Hamiltonian $H$, and energy eigenstates $|n\rangle$ with energies $E_{n}$. The thermofield-double state
@@
    is defined on two copies of the system, where $Z_{\beta}$ is the partition function. The reduced theory on each copy is in a canonical, finite-temperature state.
@@
    Universal gate set: A set of quantum gates that can be used to approximately effect every unitary operator arbitrarily precisely.